The admission fee at an amusement park is $2.75 for children and $4.80 for adults. On a certain day, 316 people entered the park, and the admission fees collected totaled 1197 dollars. How many children and how many adults were admitted?
Answer by ankor@dixie-net.com(22740)

The admission fee at an amusement park is $2.75 for children and $4.80 for adults.
On a certain day, 316 people entered the park, and the admission fees collected totaled 1197 dollars.
How many children and how many adults were admitted?
:
let a = no. of adults
let c = no. of children
:
Two equations
a + c = 316
arrange for substitution
c = -a + 316
:
4.80a + 2.75c = 1197
substitute (-a+316) for c
4.80a + 2.75(-a+316) = 1197
4.80a - 2.75a + 869 = 1197
2.05a = 1197 - 869
2.05a = 328
a = 328/2.05
a = 160 adults
and
c = 316 - 160
c = 156 children
:
:
Check this, find the actual amt from each group
4.80*160 = $768
2.75*156 = $429
----------------
total amt: 1197